SUBMILLIMETER WAVE SPECTRA OF NCS

A. MAEDA, H. HABARA^{a}, \underline{T. AMANO}, Institute for Astrophysics and Planetary Sciences, Ibaraki University, 2-1-1 Bunkyo, Mito 310-8512, Japan.

Rotational spectra of NCS in the excited bending vibrational states have not been analyzed in detail, because of difficulty in analysis due to large Renner - Teller effect and anharmonic interactions. We have extended the observation of the pure rotational spectra of NCS in the v_{2} = 1 as well as in the X^{2}Π ground state up to J = 53.5 in the submillimeter wave region. For the ground state, the analysis was straightforward. The measured transition frequencies were fit to a standard effective Hamiltonian for ^{2}Π vibronic states, and the improved molecular constants including the higher order centrifugal distortion constants were obtained. The v_{2} = 1 vibronic state splits into four vibronic sub-states, ^{2}Δ_{5/2}, ^{2}Δ_{3/2}, μ^{2}Σ and κ^{2}Σ. As a first attempt, the Δ and Σ states were fitted separately to effective Hamiltonians. The least square fittings converged by including various effective parameters concerned with the P -type doubling. However, the physical significance of these higher order parameters is not clear. It is found that the spin rotation coupling constant, γ, and the Λ-type doubling constants, p, q, are significantly different from those for the ground state. We have attempted to analyze the Δ and Σ states simultaneously by taking into account the perturbations from higher vibrational states. An isotopic species, NC^{34}S, in the X^{2}Π_{3/2} state has also been measured and r_{0} structure is derived to be r_{0}(NC)= 1.1805(1) Åand r_{0}(CS)= 1.63212(7) Å.

aNASDA postdoctral fellow